Explain the time value of money in your own words. Provide at least two examples of how you can utilize time value of money in your current occupation or in your future plans. How would you go about assessing these examples?
Answer #1

Answer

Time value of money is concept that in present the value of money is more worth than value of money in future. So early you got money (same amount in future) is more worth. This is because we can invest that money and also due to the purchasing power. In future due to inflation we can not buy the same product we can buy in present with particular amount of money.

Examples of time value of money.

Example in current occupation :

Suppose my boss say me that you wiil get $100000 salary after 5 year. And $17500 every year and interest rate is 10%.

Now which option should I prefer?

If I will take the 17500 then at the end of 5 year I will get only 87500. Which is less than 100000.

But if I will consider time value of money than I will choose option B.

How?

If I take all the year $17500 I can invest it at 10%.

So the FV of this amount is as below.

Fv=A(FVIFA n year, r %.)

=17500(6.1050).......from the fv table.

=106837.5

Here the present value of 100000 is

Pv=fv(PVIF n year, r%)

=100000(0.6209)....... From the table of PV.

Note: n=5 years and r=10%.

Example of future plan:

Now suppose I want to make house and I need 100000 to make the house. I want to make house after 5 year and interest rate is 10%. So how much should I save every year to collect 100000.

We can calculate it with the following formula

Fv=A(FVIFA n year, r %)

100000=A(6.1051).......from the tabe of annuity.

100000 /6.1051=A

16380=A

So I have to save every year 16780 and invest it with 10%.